What is the pH of a buffer system with 1.3 M C6H5ONa and 1.2 M C6H5OH (Ka= 1.0 x 10-10)

(A) 7.00

(B) 10.00

(C) 9.97

(D) 10.03

Work:
1
Expert's answer
2017-04-18T01:37:05-0400
This buffer solution is a mixture of a weak acid C6H5OH and a salt of this acid - C6H5ONa.

The pH value of basic buffer system we can calculate, using Henderson - Hasselbalch equation:

pH = pKa + lg, where pKa = - lg Ka (Ka is acid constant dissociation).

Ka(C6H5OH) = 1.0 · 10-10

We can determine the pKa value:

pKa = - lg 1.0 · 10-10 = 10

So, plug in values to find

pH = 10 + lg = 10.03

Answer: D.